# Breast Histopathology Image Retrieval by Attention-based Adversarially Regularized Variational Graph Autoencoder with Contrastive Learning-Based Feature Extraction

### v2

@@ 3.3 Graph Construction
GNNs are mostly applied on graph-based datasets. These datasets typically include nodes, node features, edges, and adjacency matrices. In contrast, non-structural datasets, such as image-based data, lack inherent graph structures. This makes graph construction an essential step for applying GNNs to such data. For the retrieval task, graph construction methods often treat subjects (e.g., images) as graph nodes ($V$), with the features of these subjects acting as node features ($X$). Edges between nodes ($E$) can be defined using distance metrics like cosine and Euclidean distances, indicating the likelihood of interaction between entities [53, 54, 55, 56]. The topology of these graphs is usually determined experimentally, employing strategies such as a pre-defined proximity threshold, a nearest neighbor rule, or a probabilistic model. To create edges between nodes, various algorithms such as $k$ approximate nearest neighbor ($k$-ANN) can be used. This approach is more time-efficient compared to the $k$-NN algorithm and is beneficial, especially in high-dimensional feature spaces, where the complexity of high dimensions slows down exact nearest neighbor searches. While there are various implementations of ANN algorithms, the ANN benchmark [57] is a widely-used library that encompasses most ANN-based methods. In our study, we utilize the fast library for approximate nearest neighbors (FLANN) to calculate distances between node embeddings, selecting the $k$ nearest nodes to each entity as adjacent nodes. The computation of graph edges is conducted as follows:
$e_{ij}=\begin{cases}1&\text{if }v_{j}\in k\text{-ANN}(v_{i})\\ 0&\text{otherwise}\end{cases}\quad(1)$
where $e_{ij}=1$ indicates the existence of an edge connection between the $i$-th and $j$-th nodes in the graph, and $e_{ij}=0$ denotes no edge. Utilizing the edges and nodes, the adjacency matrix ($A$) can be constructed where the values of the matrix ($a_{ij}$) act as indicators, quantifying of the connections between nodes $V_{i}$ and $V_{j}$.
It is worth mentioning that we use transductive learning, where the entire dataset (training, validation, and test data) is observed to build the graph. The summarized workflow to build the graph in our study is shown in Figure 3.
Figure 3: The framework of graph construction. CCL: Cluster guided contrastive learning, $K$-ANN: $K$ approximate nearest neighbors, FLANN: fast library for approximate nearest neighbors
@@ 3.4 A-ARVGAE
The overall architecture of the A-ARVGAE model is depicted in Figure 4. The proposed architecture comprises four main components: the encoder (yellow block), which also serves as the generator for adversarial training; the embedding latent space (pink block); the decoder (blue block); and the discriminator (green block), which is another essential component for adversarial training.
The constructed graph, as described in the previous section, is fed into the encoder of the A-ARVGAE model. The encoder includes a graph attention layer that performs attention operations, batch normalization, leaky ReLU, and dropout, followed by two graph convolutional layers. The attention-based convolutional operation integrating attention mechanisms into convolutional layers assigns varying weights to the neighbors, enabling the model to concentrate on the most relevant parts of the graph when learning node representations [45]. The attention mechanism aggregates information from adjacent nodes with an adaptive weight allocation to neighboring nodes. As indicated by prior studies, the use of a multi-head attention layer has been shown to yield superior performance [25] and hence we also employ two-head attention layer in our work.
Inspired by the VGAE [19], the encoder in our architecture does not encode an input into a fixed latent representation. Instead, the outputs of the graph convolutional layers are parameters (mean ($\mu$) and variance ($\sigma$)) that define a probability distribution over the possible latent representations, denoted as $Z$, characterized by a normal distribution $N(\mu,\sigma)$. Sampling from the latent space involves drawing samples from the distribution $N(\mu,\sigma)$ to produce latent node embeddings. This process employs the reparameterization trick, as introduced in [19], which enables the backpropagation of gradients through the stochastic sampling process for gradient descent optimization.
@@
The complete model integrates three loss components: a regression-based reconstruction loss (related to VGAE), a Kullback–Leibler divergence loss (related to VGAE), and a binary cross-entropy loss (related to discriminator in adversarial training). Further details regarding the model architecture and cost functions can be found in the respective studies [22, 45].
We compare the results of our employed GNN model (A-ARVGAE) with other GNN-based models, namely GAE, VGAE, and ARVGA, and also with another non-GNN-based state-of-the-art approach, called RetCCL, represented in [27]. RetCCL is a model that is recently developed for the histological image retrieval task, which is based on CCL feature extraction and nearest neighbor search to find similar images in a given dataset. The results of these comparisons are reported in Section 4 (Table 4).
Figure 4: Architecture of attention-based adversarially regularized variational graph autoencoder (A-ARVGAE). For interpretation of the references to color in this figure legend, the reader is referred to the web version of this article.
@@ 3.6 Image Embedding Database
The embeddings generated by the trained A-ARVGAE encoder in the previous step for each training image are stored in a database. This database is then utilized in the testing phase to identify the top-$K$ similar images for a given query image as explained in Section 3.7
@@ 3.7 Testing phase
As depicted in the lower part of Figure 1, for a query image, the model aims to identify the top $K$ similar images in the database. Similar to the process with training images, features are first extracted from the test image using the CCL feature extractor. These features are then fed into the trained A-ARVGAE model to obtain the embedding in a lower-dimensional space, specifically from the encoder part of the model. After acquiring the embedding of the query image, it is compared with the embeddings in the image database to identify similar images. To find the most similar images, the ranking algorithm evaluates and ranks the candidate images based on Euclidean distance. These selected images are then arranged and ranked in ascending order according to Euclidean distances. Once ranked, the top-$K$ similar images are chosen as the final output of the retrieval model.
To evaluate the performance of the developed approach, we use two evaluation metrics, namely, mean Average Precision (mAP) [58] and mean Majority Vote (mMV) [13].
mAP is the average of the precision values computed at the ranks of relevant images in the retrieved results. It is a widely used metric for evaluating the effectiveness of a retrieval system. The formula for mAP is as follows:
$mAP(k)=\frac{1}{|Q|}\sum_{q=1}^{|Q|}\frac{1}{|R_{q}|}\sum_{k=1}^{|R_{q}|}\text{Precision}(k)\cdot\text{rel}(k)\quad(17)$
Where $|Q|$ is the number of queries, $R_{q}$ is the number of relevant images for query image $q$, $\text{Precision}(k)$ is the precision at rank $k$ (the ratio of relevant images retrieved in the top $k$ results to the total number of retrieved images up to rank $k$), $\text{rel}(k)$ is an indicator function that equals 1 if the image at rank $k$ is relevant and 0 otherwise. ||$\cdot$|| ⋅ | denotes the length of the set.
mMV evaluates the frequency with which the primary diagnosis among the top-$k$ results aligns with the queried diagnosis as follows:
$mMV(k)=\frac{1}{|Q|}\sum_{q=1}^{|Q|}1\{L_{q}\in MV(\text{rel}(k))\}\quad(18)$
Where $L_{q}$ represents the ground truth diagnosis of the image, and MV is the predicted diagnosis of an image taken from the majority vote of top-$k$ samples. Similar to mAP, $\text{rel}(k)$ is an indicator function that equals 1 if the image at rank $k$ is relevant and 0 otherwise. Here, ||$\cdot$|| ⋅ | also denotes the length of the set.
@@ 4.1 Implementation Details
We used the PyTorch deep learning framework (version 2.2.0) and the PyTorch Geometric library (version 2.4.0) to implement the models. Adam optimizer was used to train the models with an initial learning rate (LR) of 0.0001. For dropout layers, a probability of 0.2 was chosen in the model architecture. All models were trained for 250 epochs, whereas in models with adversarial components, the discriminator was trained for five iterations in each epoch. To extract features from the training and test images, all images were down-sampled to 256$\times$256 pixels to fit into the utilized GPU memory. For each experiment, an identical training set (70%), validation set (10%), and test set (20%) were used to ensure fair comparisons between different experiments. The data splitting was done randomly by keeping the ratio between different classes in the training, validation and test sets. In the process of graph construction, the parameter $k$ was set to 25 for the BreakHis dataset and 15 for the BACH dataset. For the attention layer in the A-ARVGAE model, a 2-head attention layer was used. All experiments were conducted on a single machine using an NVIDIA GTX 1060 GPU and an Intel Core™ i7-9750H CPU.
@@ 4.2 Experiments
We performed two sets of experiments for each dataset in this study. In the first set, we investigated the effect of using the CCL feature extractor against other pre-trained models for the breast cancer histology image retrieval task. The results of this experiment for the BreakHis and BACH datasets are reported in Table 2 and Table 3, respectively. For a fair comparison, we used the A-ARVGAE model in all experiments.
Table 2: Evaluation of the retrieval performance for the BreakHis dataset with different feature extractors. In all experiments attention-based adversarially regularized variational graph autoencoder is used as the retrieval model.
@@
Table 3: Evaluation of the retrieval performance for the BACH dataset with different feature extractors. In all experiments attention-based adversarially regularized variational graph autoencoder is used as the retrieval model.
@@
As observed from Table 2 for the BreakHis dataset, the CCL feature extractor outperforms all other pre-trained feature extractor models across all metrics, except for 400$\times$ magnification, where DenseNet121 delivered better retrieval performance. However, for the average performance across all magnifications, CCL delivers the best results compared with all other pre-trained models, including DenseNet121. The results in Table 3 also indicate the superior performance of the CCL feature extractor compared with other pre-trained models for the BACH dataset.
Overall, the results reported in Table 2 and Table 3 for the BreakHis and BACH datasets show that the feature extractor trained on histological images outperforms feature extractors trained with non-medical images.
In the second experiment, we compare the performance of the employed A-ARVGAE model with other GNN-based and non-GNN-based models for both BreakHis and BACH datasets. The results of these experiments are presented in Table 4.
Table 4: Comparison of different models (GNN-based models including GAE, VGAE, and ARVGA, non-GNN-based model, and ours) for the BreakHis and BACH datasets for the breast cancer histological image retrieval task. GAE: graph autoencoder, VGAE: variational graph autoencoder, ARVGA: adversarial regularized variational graph autoencoder, RetCCL: retrieval with cluster-guided contrastive learning
@@
BreakHis (40$\times$) GAE 0.901 0.773
VGAE 0.962 0.934
ARVGA 0.969 0.909
RetCCL 0.983 0.9789
Ours 0.978 0.9499
BreakHis (100$\times$) GAE 0.910 0.792
VGAE 0.966 0.903
ARVGA 0.949 0.909
RetCCL 0.970 0.951
Ours 0.980 0.961
BreakHis (200$\times$) GAE 0.926 0.825
VGAE 0.955 0.900
ARVGA 0.949 0.879
RetCCL 0.920 0.920
Ours 0.964 0.930
BreakHis (400$\times$) GAE 0.924 0.828
VGAE 0.923 0.806
ARVGA 0.932 0.834
RetCCL 0.897 0.889
Ours 0.940 0.839
BreakHis Avg. GAE 0.915 0.804
VGAE 0.951 0.885
ARVGA 0.949 0.882
RetCCL 0.942 0.934
Ours 0.965 0.919
BACH (200$\times$) GAE 0.919 0.740
VGAE 0.937 0.798
ARVGA 0.935 0.817
RetCCL 0.901 0.8942
Ours 0.947 0.913
As shown in Table 4, the A-ARVGAE model surpasses all GNN-based models i.e., GAE, VGAE, and ARVGA for both datasets. The superior performance of the proposed model over ARVGA demonstrates the significance of the attention component in the employed architecture. An attention component allows the model to focus on specific parts of input data to calculate the model weights.
Compared with the RetCCL model, our approach delivers superior performance for both evaluation indices on the BACH dataset. However, for the BreakHis dataset, RetCLL shows slightly better performance for some magnifications. Nevertheless, our approach still yields superior results in other magnifications. The average results across all magnifications indicate the superior performance of our approach compared with RetCCL, based on the mAP(5) evaluation index (2.3% improved performance). However, for the MVP(5) score, RetCCL exhibits a better average performance with a 1.5% margin.
Overall, while our approach has demonstrated superior or very competitive performance in all experiments, it should be noted that GNN training typically requires large datasets [59]. The utilized model in this study was trained on the BreakHis and BACH datasets, which are relatively small.
For visual analysis, we randomly select five query images from the BreakHis and BACH test sets (four from the BreakHis with different magnifications and one from the BACH dataset) to visualize the model output (top five ranks). Figure 5 depicts the query images and model output. As the results show, the model is capable of providing relevant images (with identical classes as the query images) for most of the samples.
@@
N N N N N N
Figure 5: Five sample queries from the BreakHis and BACH datasets and their top similar retrieved images. Successful and unsuccessful retrieval are denoted by () and (), respectively. M: Malignant class, BN: Benign class, N: Normal class, BH: BreakHis
As suggested in [60], we also compare the embedding vectors of some example image pairs with different degrees of similarity in Figure 6. As shown in the figure, the embedding vectors for similar image pairs are very close to each other. In contrast, the embedding vectors for dissimilar pairs are separable.
@@ 5 Conclusion & Future Works
Recently, there has been a growing trend in utilizing GNNs to address challenges in analyzing histopathology images. This research introduces a new GNN-based approach specifically designed for breast histopathology image retrieval. We incorporated both attention mechanisms and adversarial training in a variational graph auto-encoder model to enhance retrieval performance. Moreover, instead of using conventional pre-trained models as the node feature extractor, we employed the CCL feature extractor, specifically pre-trained on histological images. Our results confirm that incorporating both attention mechanisms and adversarial training consistently exhibits superior performance compared to other GNN variants. Additionally, using the CCL feature extractor further boosts performance in our experiments. While our approach delivers promising results for breast histology image retrieval, certain aspects can be considered for further investigation in future studies.
Transformers-based models have shown impressive performance in various computer vision tasks. Some Transformer-based GNNs, such as graph transformer networks [61], have been developed for tasks such as node classification. Such models can be adapted and employed for the medical image retrieval task as well.
While we use FLANN to build the graph in our model, other state-of-the-art ANN methods and libraries, such as Faiss library  [62] or HNSW method  [8] can be exploited to enhance performance. However, it should be noted that most of these methods are slower in comparison to FLANN  [36].
@@
Finally, using well-known boosting techniques, such as test-time augmentation or ensembling, has been shown to enhance performance for tasks such as medical image classification or segmentation [63, 64, 65]. Such techniques can also be incorporated into the workflow, and their impact on the retrieval performance can be investigated.

### v3

@@ 3.3 Graph Construction
GNNs are mainly applied to graph-based datasets. These datasets typically include nodes, node features, edges, and adjacency matrices. In contrast, non-structural datasets, such as image-based data, lack inherent graph structures. This makes graph construction an essential step for applying GNNs to such data. For the retrieval task, graph construction methods often treat subjects (e.g., images) as graph nodes, with the features of these subjects acting as node features. Edges between nodes can be defined using distance metrics like cosine and Euclidean distances, indicating the likelihood of interaction between entities [50, 51]. By utilizing the edges and nodes, the adjacency matrix can be constructed where the values of the matrix act as indicators, quantifying the connections between nodes.
The topology of the graphs is usually determined experimentally, employing strategies such as a pre-defined proximity threshold, a nearest neighbor rule, or a probabilistic model. Various algorithms, such as $k$ approximate nearest neighbor ($k$-ANN), can create edges between nodes. This approach is more time-efficient than the $k$-NN algorithm and beneficial, especially in high-dimensional feature spaces, where the complexity of high dimensions slows down exact nearest neighbor searches. While there are various implementations of ANN algorithms, the ANN benchmark [52] is a widely used library encompassing most ANN-based methods. In this study, we utilize the fast library for approximate nearest neighbors (FLANN) to calculate distances between node embeddings, selecting the $k$-nearest nodes to each entity as adjacent nodes.
We use transductive learning, where the entire dataset (training, validation, and test data) is observed to build the graph. The summarized workflow of building the graph is shown in Fig. 3.
Figure 3: The framework of graph construction. $K$-ANN: $K$ approximate nearest neighbors, FLANN: fast library for approximate nearest neighbors
@@ 3.4 A-ARVGAE
The overall architecture of the A-ARVGAE model is depicted in Fig. 4. The proposed architecture comprises four main components: the encoder (yellow block), which also serves as the generator for adversarial training, the embedding latent space (pink block), the decoder (blue block), and the discriminator (green block), which is another essential component for adversarial training.
The constructed graph, as described in the previous section, is fed into the encoder of the A-ARVGAE model. The encoder includes a graph attention layer that performs attention operations, batch normalization, leaky ReLU, and dropout, followed by two graph convolutional layers. The attention-based convolutional operation integrates attention mechanisms into convolutional layers by assigning varying weights to the neighbors and enabling the model to concentrate on the most relevant parts of the graph when learning node representations [53]. As indicated by prior studies, a multi-head attention layer has yielded superior performance [17]; hence, we also employ the two-head attention layer in our work.
Inspired by the VGAE [20], the encoder in our architecture does not encode an input into a fixed latent representation. Instead, the outputs of the graph convolutional layers are parameters (mean ($\mu$) and variance ($\sigma$)) that define a probability distribution over the possible latent representations, denoted as $Z$, characterized by the Normal distribution $N(\mu,\sigma)$. Sampling from the latent space involves drawing samples from the distribution $N(\mu,\sigma)$ to produce latent node embeddings. This process employs the reparameterization trick, as introduced in [20], which enables the backpropagation of gradients through the stochastic sampling process for gradient descent optimization.
@@
The complete model integrates three loss components: a regression-based reconstruction loss (associated with VGAE), a Kullback–Leibler divergence loss (associated with VGAE), and a binary cross-entropy loss (related to discriminator in adversarial training). Further details regarding the model architecture and cost functions can be found in the respective studies [14, 53].
As an ablation study, we compare the results of our employed GNN model (A-ARVGAE) with other GNN-based models, namely GAE, VGAE, and ARVGA. The results of these comparisons are reported in Section 4 (Table 4).
Figure 4: The Architecture of attention-based adversarially regularized variational graph autoencoder (A-ARVGAE). For interpretation of the references to color in this figure legend, the reader is referred to the web version of this article.
@@ 3.6 Image Embedding Database
The embeddings generated by the trained A-ARVGAE encoder in the previous step for each training image are stored in a database. This database is then utilized in the testing phase to identify the top-$K$ similar images for a given query image as explained in Section 3.7.
@@ 3.7 Testing phase
As depicted in the lower part of Fig. 1, the model aims to identify similar top-$K$ images in the database for a query image. Similar to the process with training images, features are first extracted from the test image using the pre-trained feature extractors. These features are then fed into the trained A-ARVGAE model to obtain the embedding in a lower-dimensional space, specifically from the encoder part of the model. After acquiring the embedding of the query image, it is compared with the embeddings in the image database to identify similar images. To find the most similar images, the ranking algorithm evaluates and ranks the candidate images based on Euclidean distance. These images are then arranged and ranked in ascending order according to Euclidean distances. Once ranked, similar top-$K$ images are chosen as the final output of the retrieval model.
To evaluate the performance of the developed approach, we use two widely-employed evaluation metrics for image retrieval, namely, mean Average Precision (mAP ($k$)) [54] and mean Majority Vote (mMV ($k$)) [35], where $k$ represents the top k results.
@@ 4.1 Implementation Details
We implemented the graph neural network using the PyTorch deep learning framework and the PyTorch Geometric library. We used the Keras framework for feature extraction with pre-trained CNN models, where the feature extraction from BioMedCLIP, UNI, and CCL models was performed using their respective repositories on Hugging Face and GitHub  [^1] https://huggingface.co/microsoft/BiomedCLIP-PubMedBERT_256-vit_base_patch16_224
https://github.com/mahmoodlab/UNI
https://github.com/Xiyue-Wang/RetCCL
Adam optimizer was used to train the models with an initial learning rate (LR) of 0.0001. For dropout layers, a probability of 0.2 was chosen in the model architecture. All models were trained for 250 epochs, but for the models with adversarial components, the discriminator was trained for five iterations in each epoch. All images were down-sampled to 256$\times$256 pixels to fit into the utilized GPU memory to extract features from the training and test images. For each experiment, an identical training set (70%), validation set (10%), and test set (20%) were used to ensure fair comparisons between different experiments. The data splitting was done randomly by keeping the ratio between different classes in the training, validation, and test sets. In the graph construction process, the parameter $k$ was set to 25 for the BreakHis dataset and 15 for the BACH dataset. For the attention layer in the A-ARVGAE model, a 2-head attention layer was used. All experiments were conducted on a single machine using an NVIDIA GTX 1060 GPU and an Intel Core™ i7-9750H CPU.
@@ 4.2 Experiments
We performed two sets of experiments for each dataset in this study. In the first set, we investigated the effect of using the medical foundation model feature extractors against other pre-trained CNN models for the histology image retrieval task for breast cancer. The results of this experiment for the BreakHis and BACH datasets are reported in Table 1, Table 2, and Table 3. For the BreakHis dataset, Table 1 shows the results per image magnification, while Table 2 provides the average results across magnifications. For a fair comparison, we used the A-ARVGAE model in all experiments.
Table 1: Evaluation of the retrieval performance for the BreakHis dataset with different feature extractors across different magnifications. The best results for convolutional neural networks and foundation models are highlighted by underlining and bold font, respectively. The attention-based adversarially regularized variational graph autoencoder is used in all experiments as the retrieval model.
@@
Table 2: Evaluation of the average retrieval performance for the BreakHis dataset with different feature extractors. The best results for convolutional neural networks and foundation models are highlighted by underlining and bold font, respectively. The attention-based adversarially regularized variational graph autoencoder is used in all experiments as the retrieval model.
Magnification Feature Extraction mAP(5) mMV(5)
@@
Table 3: Evaluation of the retrieval performance for the BACH dataset with different feature extractors. The best results for convolutional neural networks and foundation models are highlighted by underlining and bold font, respectively. The attention-based adversarially regularized variational graph autoencoder is used in all experiments as the retrieval model.
@@
As observed from Table 1, for the BreakHis dataset, models trained with foundation model features outperform all models trained with CNN features across both metrics, except for 400$\times$ magnification, where DenseNet121 delivered better retrieval performance. As shown in Table 2, for the average performance across all magnifications, the employed medical foundation models deliver superior performances compared with all other CNN models, including DenseNet121. Based on the average results for the BreakHis dataset, the top-performing foundation model achieves a 1.7% and 4.5% improvement over the best CNN model in terms of mAP(5) and mMV(5), respectively.
The results in Table 3 also indicate the superior performance of the medical foundation model feature extractors compared with other CNN models for the BACH dataset. Based on the results for the BACH dataset, the top-performing foundation model achieves a 7.7% and 15.5% improvement over the best CNN model in terms of mAP(5) and mMV(5), respectively.
Overall, the results presented in Table 2 and Table 3 for the BreakHis and BACH datasets indicate that feature extractors trained on medical/histological images outperform those trained on non-medical images. Additionally, comparing foundation models shows highly competitive performance across all foundation models on the BreakHis dataset, while UNI achieves the best performance on the BACH dataset.
In the second experiment, we compare the performance of the employed A-ARVGAE model with other GNN-based models for both BreakHis and BACH datasets. The results of these experiments are presented in Table 4. The UNI model was used as the graph feature extractor to derive the results in this table.
Table 4: Comparison of different GNN-based models for the BreakHis and BACH datasets for the breast cancer histological image retrieval task. The UNI model was used as the graph feature extractor in all experiments. GAE: graph autoencoder, VGAE: variational graph autoencoder, ARVGA: adversarial regularized variational graph autoencoder, A-ARVGAE: attention-based adversarially regularized variational graph autoencoder.
@@
BreakHis (40$\times$) GAE 0.897 0.783
VGAE 0.960 0.894
ARVGA 0.930 0.817
A-ARVGAE 0.985 0.954
BreakHis (100$\times$) GAE 0.896 0.768
VGAE 0.960 0.937
ARVGA 0.949 0.922
A-ARVGAE 0.973 0.946
BreakHis (200$\times$) GAE 0.907 0.771
VGAE 0.930 0.855
ARVGA 0.938 0.865
A-ARVGAE 0.975 0.910
BreakHis (400$\times$) GAE 0.925 0.784
VGAE 0.934 0.790
ARVGA 0.930 0.817
A-ARVGAE 0.938 0.850
BreakHis Avg. GAE 0.906 0.776
VGAE 0.946 0.869
ARVGA 0.936 0.855
A-ARVGAE 0.967 0.915
BACH (200$\times$) GAE 0.941 0.884
VGAE 0.962 0.865
ARVGA 0.962 0.855
A-ARVGAE 0.976 0.942
As shown in Table 4, the A-ARVGAE model surpasses all GNN-based models, i.e., GAE, VGAE, and ARVGA of both datasets. The superior performance of the proposed model over ARVGA demonstrates the significance of the attention component in the employed architecture. An attention component allows the model to focus on specific parts of input data to calculate the model weights. The average results indicate that the A-ARVGAE model achieves 2.1%/4.6% and 1.4%/5.8% higher performance in terms of mAP and mMV compared to the second-best GNN model for the BreakHis and BACH datasets.
For visual analysis, we randomly select five query images from the BreakHis and BACH test sets (four from the BreakHis with different magnifications and one from the BACH dataset) to visualize the model output (top five ranks). Fig. 5 depicts the query images and model output, where the model (A-ARVGAE with UNI feature extractor) provides relevant images (with identical classes acting as the query images) for most samples.
@@
BN BN BN BN BN BN
Figure 5: Five sample queries from the BreakHis and BACH datasets and their top similar retrieved images. Successful and unsuccessful retrieval are denoted by the symbols () and (), respectively. M: Malignant class, BN: Benign class, BH: BreakHis
As suggested in [55], we also compare the embedding vectors of some example image pairs with different degrees of similarity in Fig. 6, in which the embedding vectors for similar image pairs are very close. In contrast, the embedding vectors for dissimilar pairs are separable.
Similar Pairs Refer to caption
Refer to caption Refer to caption
Dissimilar Pairs Refer to caption
Refer to caption Refer to caption
Similar Pairs Refer to caption
Refer to caption Refer to caption
Dissimilar Pairs Refer to caption
Refer to caption Refer to caption
@@ 5 Conclusion & Future Works
Recently, there has been a growing trend in utilizing GNNs to address challenges in analyzing histopathology images. This research introduces a new GNN-based approach specifically designed for breast histopathology image retrieval. We incorporated attention mechanisms and adversarial training in a variational graph auto-encoder model to enhance retrieval performance. Moreover, instead of using conventional pre-trained CNN models as the node feature extractor, we employed the features from medical foundation models, specifically pre-trained on medical/histological images. Our results confirm incorporating attention mechanisms and adversarial training outperforms the other GNN variants. Additionally, using the feature extracted from foundation models further boosts performance in our experiments. While our approach delivers promising results for breast histology image retrieval, certain aspects can be considered for further investigation in future studies.
Transformers-based models have shown impressive performance in various computer vision tasks. Some Transformer-based GNNs, such as graph transformer networks [56], have been developed for tasks such as node classification. Such models can be adapted and employed for medical image retrieval.
While we use FLANN to build the graph in our model, other state-of-the-art ANN methods and libraries, such as the Faiss library [57] or the HNSW method [8], can be exploited to enhance performance. However, most of these methods are slower than FLANN  [33].
@@
We used three medical foundation models in the study as the graph node feature extractors. However, alternative foundation models could also be explored by considering recent advancements in this field and the availability of other medical and non-medical foundation models. We selected these specific medical foundation models because they have demonstrated superior performance than several other foundation models across various computer vision tasks in former studies  [21, 23, 58].
Finally, well-known boosting techniques, such as test-time augmentation or ensembling, can enhance performance for medical image classification or segmentation  [59, 60, 61]. The resulting retrieval performance can be investigated by incorporating such techniques into the workflow.